OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL! Suppose in a proof of a summation formula 1+5+25+...+5^k -1=1/4 (5^k -1) by mathematical induction, you show the formula valid for L=1 and assume that it is valid for L=k. What is the next equation in the induction step of this proof? a 1+5+25+...+5^k-1 +5^k+1-1 = 1/4(5^k -1) +1/4(5^k+1 -1) b 1+5+25+...+5^k +5^k+1 = 1/4(5^k -1)+5^k+1-1 c 1+5+25+...+5^k-1 +5^k+1-1 =1/4 (5^k -1) + 5^k+1-1 d 1+5+25+...+5^k-1 +5^k+1-1 = 1/4 (5^k+1 -1) + 5^k+1-1

OpenStudy (anonymous):

add the next kth term to each side maybe?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

well, you need to know if its true for (k+1) terms, and you know its true for (k) terms soo add the next term to each side ....

OpenStudy (anonymous):

in other words: P(k) = format(k) P(k) + (k+1) = format(k) + (k+1)

OpenStudy (anonymous):

so would it be B because it has 5^k + 5^k+1 ?

OpenStudy (anonymous):

it hard to tell, but i think B is good on the right, but bad on the left

OpenStudy (anonymous):

1+5+25+...+ 5^k-1 = 1/4 (5^k-1) ^^^^^ let k = k+1 1+5+25+...+ 5^k-1 + 5^k+1-1 = 1/4 (5^k-1) + 5^k+1-1

OpenStudy (anonymous):

I see! Thank you so much!!

OpenStudy (anonymous):

youre welcome